On a cold winter day, the outside temperature is -20°C and the inside temperature is maintained at 20°C. There is a net heat flow to the outside through the walls, roof, etc., of 25 kW. What is the rate of increase of the entropy of the universe as a result of this heat flow?
Answer #1

Okay,

so

Shot = 25000 / 293 = 85.3 W/K

Scold = 25000 / 253 = 98.8 W/K

Therefore,

rate of increase of the entropy of the universe

S = 98.8 - 85.3

S = 13 W/K